OpenStudy (anonymous):

Is this answer correct: Question: simplify the expression: cos(270degrees + x) Answer: 0.762

OpenStudy (experimentx):

I don't think so ..

OpenStudy (anonymous):

Dont I use the sum difference formula?

OpenStudy (anonymous):

No. There is a variable so it cannot be as the answer has to depend on x

OpenStudy (experimentx):

I'll tell you one thing ... if you ever got values greater than 1 from Sin or Cos ... then you will always be wrong.

OpenStudy (anonymous):

this is what I did, cos (theta + beta) = cos 270 cos x + sin270 sinx

OpenStudy (anonymous):

then I get 0.975 + 0.2297

OpenStudy (anonymous):

wait, I added something differently

OpenStudy (experimentx):

cos(270+x) = - sin(x)

OpenStudy (anonymous):

shouldnt the answer be 1.20?

OpenStudy (anonymous):

OHHH! I see where I went wrong

OpenStudy (anonymous):

I added intead of substracting

OpenStudy (anonymous):

Shouldt the answer be 0.7462?

OpenStudy (anonymous):

try this... use your graphing calculator (DEGREE mode, so change your x,y scale): y1 = cos(270 + x) y2 = -sin(x) look at their graphs...

OpenStudy (experimentx):

sorry, sin(x) must be your answer ...stupid me

OpenStudy (anonymous):

Im confused, -sinx or sinx?

OpenStudy (phi):

oops cos(270degrees + x)= cos(270)cos(x)-sin(270)sin(x)= +sin(x)

OpenStudy (anonymous):

oh boy:)

OpenStudy (anonymous):

ok. y2 = sinx just going with the flow...

OpenStudy (phi):

I take back saying Open did it correctly: this is what I did, cos (theta + beta) = cos 270 cos x + sin270 sinx should be cos 270 cos x - sin270 sinx

OpenStudy (anonymous):

Thats what i said after:) I added instead of substract

OpenStudy (phi):

now if you have a value for x, you can go further.

OpenStudy (anonymous):

what they're trying to say is... you shouldn't be getting a numeric answer.. cos(270 + x) = sinx is an identity...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

try it out on your graphing calculator....

OpenStudy (anonymous):

thats weird, when I graph it, I get like a line on opposite ends, and nothin in the middle

OpenStudy (anonymous):

degree mode..

OpenStudy (anonymous):

Yeah, It is on degree

OpenStudy (anonymous):

window [-360, 360] for the x [-2,2] for the y

OpenStudy (anonymous):

ok, not I see it. I did something wrong with the zoom

OpenStudy (anonymous):

I get waves in

OpenStudy (anonymous):

I get waves in my graph

OpenStudy (anonymous):

you only have equations in y1 and y2 only? nothing in the others?

OpenStudy (anonymous):

yes, y1 = cos(270+x) y2 = -sin(x)

OpenStudy (anonymous):

y2 = sin(x)

OpenStudy (anonymous):

Oh sorry, I tohught you said -sin(x) before, Ok Im chaging it

OpenStudy (anonymous):

i did but i wsa just going with experinent and phi said....

OpenStudy (anonymous):

So now I get like a wave starting in positive side and going up and down. Is this correct?

OpenStudy (anonymous):

So basically, y1 and y2, will generate the same graph right?

OpenStudy (anonymous):

Thats why it is an identity

OpenStudy (anonymous):

the graph in y1 should be exactly on y2... that shows this is an identity.

OpenStudy (anonymous):

yep

OpenStudy (phi):

Here's what it should look like http://www.wolframalpha.com/input/?i=plot%28cos%283pi%2F2%2Bx%29%29%2C+sin%28x%29

OpenStudy (anonymous):

yeah thats the graph I got:)

OpenStudy (anonymous):

Ok so backing it up a bit, when using the sum difference formulae

OpenStudy (anonymous):

but look at what phi did on his earlier post... he actually proved cos(270+x) = sinx

OpenStudy (anonymous):

How do I get that exactly?

OpenStudy (anonymous):

I cant really put it in py calculator cause il get decimal answer

OpenStudy (phi):

cos (270 + x) = cos 270 cos x - sin270 sinx now find cos(270)=0 sin(270)=-1 so you get 0 - -sin(x) =sin(x) as your answer

OpenStudy (anonymous):

Okok. that makes sense. Thank you so much guys:)

OpenStudy (phi):

How were you getting a number out of this?